# Cislunar Space Situational Awareness Constellation Design and Planning with Facility Location Problem

### v3

@@ 6.2.2 Low-Energy Transfer Transit Window
Figure 14 shows the sensor steering history for architectures in Figure 13. In all cases, all observers are located on one side of the region of interest; as a result, the azimuth angle remains mostly between $-30^{\circ}$ and $30^{\circ}$. With the $p=2$ case utilizing Southern halos, the elevation angle across observers and across time is skewed on the positive side, while with the $p=3,4,5$ cases utilizing Northern halos, the elevation angle is skewed on the negative side. Figure 15 shows the sensor pointing directions for $\bar{m}_{\rm crit}=20$ and $p=4$ from Figure 13(c). As the Northern halo has a larger out-of-plane amplitude than the region of interest, sensors are pointed downward around the apolunes along the Northern halo. We also note that over the first 10 time steps, the steering directions of each observer have slight variations from one another due to the variation in their respective phase along the LPO.
@@ 7 Conclusions
In this work, the TE-$p$-MP has been proposed for CSSA. The formulation consists of determining the $p$ optimal observer locations along with the orientation of their optical sensors over time to observe a given set of static or dynamic targets. The TE-$p$-MP, inspired from the classical $p$-Median problem, is a BLP discretized in time, candidate locations of observer spacecraft, and sensor pointing directions. This formulation simultaneously solves for the constellation design and the sensor-tasking problems. The formulation has been developed together with a custom LM algorithm, based on the relaxation of complicating constraints and custom heuristics leveraging the problem structure as well as the geometry of the candidate facility locations. A solution to the TE-$p$-MP is a CSSA constellation together with sensor-tasking for each observer that observes a prescribed set of targets.
@@ 7.1 Parameters of Synodic-Resonant LPOs
Family $M$:$N$ $x_{0}$ $z_{0}$ $\dot{y}_{0}$ Period, TU Stability Slots $b$
@@
[…]
@@ 7.2 LPOs used in TE-$p$-MP
[…]
@@
halo L2N (9:2) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (1:1) $\times$ 1
@@
lyapunov L1 (1:1) $\times$ 2
lyapunov L1 (1:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
lyapunov L2 (3:2) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
lyapunov L1 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 1
@@
[…]
@@
halo L2N (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (1:1) $\times$ 2
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
lyapunov L1 (1:1) $\times$ 3
@@
lyapunov L2 (1:1) $\times$ 3
lyapunov L1 (1:1) $\times$ 3
@@
[…]
@@
halo L2N (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyN (1:1) $\times$ 1
butterflyS (1:1) $\times$ 2
butterflyN (1:1) $\times$ 2
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
lyapunov L1 (1:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
[…]
@@
halo L2N (4:1) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (3:2) $\times$ 1
butterflyN (1:1) $\times$ 2
lyapunov L1 (9:4) $\times$ 1
butterflyN (1:1) $\times$ 4
@@
halo L2N (9:4) $\times$ 1
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
lyapunov L1 (1:1) $\times$ 3
lyapunov L2 (1:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 3
lyapunov L1 (3:2) $\times$ 1
@@
[…]
@@
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 1
halo L2S (2:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 2
halo L2N (3:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 2
halo L2N (3:1) $\times$ 2
@@
[…]
@@
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 2
halo L2S (2:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 3
@@
[…]
@@
halo L2N (2:1) $\times$ 2
halo L2S (2:1) $\times$ 2
halo L2S (2:1) $\times$ 4
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
[…]
@@
halo L2N (2:1) $\times$ 2
halo L2S (2:1) $\times$ 3
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 4
@@
halo L2N (9:4) $\times$ 1
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
lyapunov L1 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 4

### v4

@@ 6.2.2 Low-Energy Transfer Transit Window
Figure 14 shows the sensor steering history for architectures in Figure 13. In all cases, all observers are located on one side of the region of interest; as a result, the azimuth angle remains mostly between $30^{\circ}$ and $30^{\circ}$. With the $p=2$ case utilizing Southern halos, the elevation angle across observers and across time is skewed on the positive side, while with the $p=3,4,5$ cases utilizing Northern halos, the elevation angle is skewed on the negative side. Figure 15 shows the sensor pointing directions for $\bar{m}_{\rm crit}=20$ and $p=4$ from Figure 13(c). As the Northern halo has a larger out-of-plane amplitude than the region of interest, sensors are pointed downward around the apolunes along the Northern halo. We also note that over the first 10 time steps, the steering directions of each observer have slight variations from one another due to the variation in their respective phase along the LPO.
@@ 7 Conclusions
In this work, the TE-$p$-MP has been proposed for CSSA. The formulation consists of determining the $p$ optimal observer locations along with the orientation of their optical sensors over time to observe a given set of static or dynamic targets. The TE-$p$-MP, inspired by the classical $p$-Median problem, is a BLP discretized in time, candidate locations of observer spacecraft, and sensor pointing directions. This formulation simultaneously solves for the constellation design and the sensor-tasking problems. The formulation has been developed together with a custom LM algorithm, based on the relaxation of complicating constraints and custom heuristics leveraging the problem structure as well as the geometry of the candidate facility locations. A solution to the TE-$p$-MP is a CSSA constellation together with sensor-tasking for each observer that observes a prescribed set of targets.
@@ 7.1 Parameters of Synodic-Resonant LPOs
Family $M$:NN $x_{0}$ $z_{0}$ $\dot{y}_{0}$ Period, TU Stability Slots $b$
@@
Number of slots $b$ is based on $\Delta t_{b}=12$  $\text{\,}\mathrm{h}\mathrm{o}\mathrm{u}\mathrm{r}\mathrm{s}$, resulting in 1212 slots in total.
@@ 7.2 LPOs used in TE-$p$-MP
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (9:2) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Btrfly. (S) (1:1) $\times$ 1
Btrfly. (N) (1:1) $\times$ 1
@@
L1 Lyap. (1:1) $\times$ 2
L1 Lyap. (1:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
L2 Lyap. (3:2) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
L1 Lyap. (1:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 1
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (9:2) $\times$ 1
Btrfly. (S) (1:1) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Btrfly. (S) (1:1) $\times$ 1
Btrfly. (N) (1:1) $\times$ 2
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
L1 Lyap. (1:1) $\times$ 3
@@
L2 Lyap. (1:1) $\times$ 3
L1 Lyap. (1:1) $\times$ 3
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (9:2) $\times$ 1
Btrfly. (S) (1:1) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Btrfly. (N) (1:1) $\times$ 1
Btrfly. (S) (1:1) $\times$ 2
Btrfly. (N) (1:1) $\times$ 2
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
L1 Lyap. (1:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 3
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (4:1) $\times$ 1
Btrfly. (S) (1:1) $\times$ 1
Btrfly. (N) (3:2) $\times$ 1
Btrfly. (N) (1:1) $\times$ 2
L1 Lyap. (9:4) $\times$ 1
Btrfly. (N) (1:1) $\times$ 4
@@
L2 Halo (N) (9:4) $\times$ 1
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
L1 Lyap. (1:1) $\times$ 3
L2 Lyap. (1:1) $\times$ 2
@@
L1 Lyap. (1:1) $\times$ 3
L1 Lyap. (3:2) $\times$ 1
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 2
@@
L1 Lyap. (1:1) $\times$ 2
L2 Halo (N) (3:1) $\times$ 2
@@
L1 Lyap. (1:1) $\times$ 2
L2 Halo (N) (3:1) $\times$ 2
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 3
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 3
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 3
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 4
@@
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
$\bar{m}_{\mathrm{crit}}$ B&B (3600  $\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}$) Lagrangian Method
@@
L2 Halo (N) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 3
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 4
@@
L2 Halo (N) (9:4) $\times$ 1
L2 Lyap. (3:2) $\times$ 1
Btrfly. (N) (2:1) $\times$ 1
L2 Lyap. (1:1) $\times$ 1
@@
L1 Lyap. (1:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 4